Find the general and singular solution of xp^2-yp+a=0
1
Expert's answer
2020-12-22T19:47:18-0500

Find the general and singular solution of

xp2-yp+a=0


Make y the subject of the formula

y=xp+a/p (Since this is a clairaut’s equation)

It general solution will be given by

put p=c

Then y=cx+ a/c

The general solution is   y=cx+ a/c

Differentiate with respect to c

0=x(1) - a/c2 

c2=a/x

c="\sqrt{a/c}"  

Put c into the general equation 

y=x"\sqrt{a/c}" +a"\sqrt{x/a}"

The singular solution is y=x"\sqrt{a/c}" +a"\sqrt{x/a}"